the sum of the ages of petra and her mother is 45. her mother is 9 years more than twice as old as petra. how old are petra and her mother
Answer by shahid(44)

let x years be the age of petra and y be the age of her mother.according to conditions
x+y=45 .......(1)
y=2x+9........(2) from this equ: y=2x+9 putting in equ:(1)
x+2x+9=45
3x=45-9
3x=36
x=12 putting this in equ:2
y=2(12)+9
y=33
hence petra's age is 12 and her mother's age is 33